OpenStudy (anonymous):

Janz is putting turf on a triangular portion of a garden around a circular fountain. If the turf costs $10 a square meter, how much will it cost Janz? Round to the nearest cent.

OpenStudy (anonymous):

$ 5611.84 $ 4480.87 $ 5432.84 $ 6480.00

OpenStudy (ja1):

How big is the truf?

OpenStudy (anonymous):



OpenStudy (anonymous):

crappy drawing sorry

OpenStudy (anonymous):

opp i figures it out... i got it right... it was 4480.87

OpenStudy (ja1):

\[2\times \pi \times r\]

OpenStudy (ja1):

That is the formula for the circumference

OpenStudy (ja1):

So let's substitute some terms: \[2\times3.14\times6\]

OpenStudy (ja1):

So 2 * 3.14 = ?

OpenStudy (ja1):

@drotzmann

OpenStudy (anonymous):

i figures it out i already knew that the area of the circle was 113.04m and theni found out the area of the triangle was 561.18m. so the area of garden was 448.14m....and you multiply that by 10 and it is around $4,480

OpenStudy (ja1):

Okay then.... xD here is a medal then